# Signal Processing over Time-Varying Graphs: A Systematic Review

### v1

@@ Abstract
As irregularly structured data representations, graphs have received a large amount of attention in recent years and have been widely applied to various real-world scenarios such as social, traffic, and energy settings. Compared to non-graph algorithms, numerous graph-based methodologies benefited from the strong power of graphs for representing high-dimensional and non-Euclidean data. In the field of Graph Signal Processing (GSP), analogies of classical signal processing concepts, such as shifting, convolution, filtering, and transformations are developed. However, many GSP techniques usually postulate the graph is static in both signal and typology. This assumption hinders the effectiveness of GSP methodologies as the assumption ignores the time-varying properties in numerous real-world systems. For example, in the traffic network, the signal on each node varies over time and contains underlying temporal correlation and patterns worthy of analysis. To tackle this challenge, more and more work are being done recently to investigate the processing of time-varying graph signals. They cope with time-varying challenges from three main directions: 1) graph time-spectral filtering, 2) multi-variate time-series forecasting, and 3) spatiotemporal graph data mining by neural networks, where non-negligible progress has been achieved. Despite the success of signal processing and learning over time-varying graphs, there is no survey to compare and conclude the current methodology for GSP and graph learning. To compensate for this, in this paper, we aim to review the development and recent progress on signal processing and learning over time-varying graphs, and compare their advantages and disadvantages from both the methodological and experimental side, to outline the challenges and potential research directions for future research.
Index Terms:
Graph Signal Processing, Time-Varying Graphs, Graph Machine Learning, Graph Neural Networks
I Introduction
Classical signal processing techniques are developed for signals defined on regular domains, such as time and space, the underlying structure of which can be represented by a uniform grid. In recent years, the processing of signals residing on non-Euclidean domains has gradually become more and more important in real-world practice, where the edges describe the relationships between the signal elements. Motivated by this, Graph Signal Processing (GSP) has arisen as a powerful framework for analyzing and processing signals over irregular domains. The application of GSP has covered many fields including transportation networks [1], temperature networks [2], and brain connectivity networks [3].
Conventional multi-variate algorithms can be applied to signals and data on graphs, as these signals also fall into the category of multi-variate signals. It is often more advantageous to use GSP-based methods when there is a topological structure associated with the data, especially in cases where the graph structure plays a crucial role in the dynamics of the underlying data. Non-graph methods often rely on independent or identically distributed assumptions, which fail to capture the relational dependencies between data points that are naturally encoded in a graph. This urges the development of algorithms and models that are designed specifically to address these limitations of applying conventional methods to graph signals.
Analogous to classical discrete signal processing, GSP extends concepts such as graph shift, graph convolution, Graph Fourier Transform (GFT), Graph Wavelet Transform (GWT), and basic filters like low-pass, high-pass, and band-pass filters. The notions of FIR and IIR filters are also extended in GSP [4, 5, 6, 7]. GSP-based algorithms are capable of solving various classical machine learning tasks, such as classification [8] and clustering [9]. GSP techniques also form the foundation of spectral graph-based deep learning algorithms, such as ChebNet [10] and Graph Convolutional Networks (GCN) [8], where nonlinear activations are combined with GSP methods to enable nonlinear modeling. The GSP components within graph neural networks provide model interpretability, which was previously lacking in non-graphical deep learning algorithms [11].
Nonetheless, the world is dynamic, which makes the signal vary over time, leading to the demand for time-varying signal processing. The signals captured in the real world are often incomplete and corrupted by noise. Let us assume that we are in a challenging missing-not-at-random scenario, where certain nodes consistently have missing data. In practice, this type of missing observation often corresponds to sensors that have failed or been removed from [12]. This scenario makes it difficult for non-graph methods to produce predictions accurately, as inferring model parameters can be challenging due to the missing observations and noise. Furthermore, these consistent missing observations also severely affect the accuracy if the attempt is to use correlation-based methods to infer the signals. Conversely, GSP can leverage the graph’s topological information to conduct spatial inference in addition to the temporal predictions. As an extension of GSP, Time-Varying Graph Signal Processing (TV-GSP) aims to analyze and process the signals that are residing on graphs and evolving with time. This extension is critical in many real-world situations. Examples include traffic flow in urban networks, the spread of information in social networks, and brain activity patterns measured over time. TV-GSP provides the means for understanding how these signals evolve and how they can be leveraged to extract meaningful information. A key challenge in TV-GSP is how to effectively model the joint variation of signals over both the spatial domain and temporal domain. One approach is to treat time as an additional dimension and to extend traditional GSP techniques to the spatiotemporal domain. This involves constructing joint spatiotemporal representations, where the graph structure and temporal dynamics are simultaneously taken into account.
In this review, we aim to provide an overview to present the recent progress in TV-GSP. Specifically, we first classify the current TV-GSP works into three categories according to their methodology. Then, we discuss their advantages and disadvantages in different settings. Finally, we propose several new potential research directions in this area. The main contributions are summarized as follows:
We provide a comprehensive review of the fundamental concepts and characteristics of TV-GSP, construct the knowledge system and systematically investigate the key techniques and models.
We present in detail the current mainstream TV-GSP algorithms and analyze in depth their advantages and disadvantages.
We discuss the challenges faced in TV-GSP highlighting potential future research directions and technological trends.
@@ I Introduction
The remainder of this paper is organized as follows. In Section II, we begin with important definitions and background knowledge about TV-GSP. Sections III, IV, V, and VI separately provide a comprehensive review and discussion of the existing TV-GSP approaches following the taxonomy demonstrated in Figure 1. Section VII summarizes the commonly used datasets, the corresponding application, and evaluation metrics. Finally, Section VIII concludes and indicates the future direction of the TV-GSP.
II Preliminaries
The advantage of using graphs to represent multi-variate signals is that graphs capture and represent the irregular interactions in data. The underlying topological structures of a graph signal are often formed to represent the pairwise relationship between two nodes through the connection of edges. Since an edge between two nodes is a pairwise relationship by definition, we can intuitively form and remove edges based on some measures of the strength of the relationship between the two data on each node. When this process is repeated on all the data pairs, a graph is formed on the global level.
II-A Definition of Time-Varying Graphs and Signals
In conventional GSP that operates on static data, the main research focus is placed on a static graph that is defined by a set of nodes and similarly a set of edges that represent the connections between nodes. In this review, the graph is always assumed as undirected unless otherwise specified, and can be either weighted or unweighted. A graph signal is defined as a multi-dimensional vector, recording the value of each node in . We further denote the number of nodes with .
@@ II-A Definition of Time-Varying Graphs and Signals
The adjacency matrix of comprises the topological information by storing the state of connectivity between nodes. To be specific, the entry of is the edge weight between nodes and . If is unweighted, indicates there is an edge between nodes and . Conversely, means there is no connection between and . For an undirected graph, the adjacency matrix is always symmetric. If is undirected and unweighted, the number of edges connected to a node is defined as the node degree , which can be represented by a diagonal matrix called the degree matrix . In the weighted case, the degree of a node is the sum of all edge weights connected to it, rather than simply counting the number of edges.
Since the real-world graph is rarely static and mostly dynamic, where the node values, edge weights, and the topology are all probably evolving over time, applying GSP techniques on every single time instance is obviously improper due to neglect of temporal dependency. Thus, the TV-GSP came into being. However, different TV-GSP algorithms are proposed for different sorts of time-varying graphs. In this review, we additionally divide the time-varying graphs into two classes from the evolutionary perspective. The first class is referred to as the Node-Varying Graphs (NVG), where the node signals vary over time while other graph elements, such as the topology or edge weights, remain static. Both time-vertex graphs, commonly used in GSP, and spatiotemporal graphs, frequently seen in GNNs, can be categorized under NVGs, as they focus on temporal variations in node signals. An illustration of NVG is shown in Fig. 2(a).
The second class is named Dynamic Graph (DG), where the varying signal not only appears at nodes but also happens in the graph topology itself. Correspondingly, the approach for processing NVG is named NVGSP, and the approach for processing dynamic graphs is named DGSP. The representative works in these two classes are shown in Fig. 1. Given multi-dimensional series data , where is the number of vertices and is the number of time points. For such data, time-varying graphs are efficient structures to characterize the relations between different vertices in a certain spatial and temporal range. A time-varying graph can be represented by the notation , where is the vertices set, and respectively denote the edges set and adjacency matrix at time t. In most time-varying graph modeling scenarios, the size of is constant, the size of can be time-varying or time-constant, and also changes with . In general, from the perspective of connectivity, graphs can be directed, undirected, weighted, and unweighted. In this review, we additionally divide the time-varying graphs into two classes from the evolutionary perspective.
II-B Graph Signal Processing
@@
The graph Laplacian matrix , which combines information from and , is defined as . The Graph Fourier Transform (GFT) is based on the eigenvector decomposition
@@
where is the orthonormal eigenvector matrix and is the diagonal matrix of eigenvalues . The GFT transforms a graph signal from the spatial domain to the spectral domain by projecting onto :
@@
The GFT of the time-varying graph signal (hourly temperature) in Fig. 2(a) is shown in Fig. 2(b). Spectral-domain operations can be performed similarly to those in the classical Fourier transform by defining a filter , which is then applied using the convolution property of the Fourier transform. A basic yet complete graph convolution procedure to apply the filter to in the spectral domain and transform it back into the spatial domain:
In GSP, concepts like low-pass, high-pass, band-pass, and band-reject filters can be analogously implemented through filter design [7]. Wavelets on graphs can be constructed by applying a kernel function to the decomposition of the graph Laplacian in (1) to achieve localized representations [6]. A graph signal is spectrally sparse when it is bandlimited in the frequency domain. For example, a bandlimiting filter is defined based on a frequency set , in which , with if and 0 otherwise [13].
The spectral graph convolution can be approximated into a spatial graph convolution by using Chebyshev polynomial approximation for a series of Chebyshev polynomial defined over the interval [14]:
@@
A graph signal with a reduced number of nodes, sampled or masked according to a node sampling set , is sparse in the spatial domain. The matrix represents the sampling or masking matrix, containing only non-zero diagonal elements given by . The matrix can be designed to introduce spatial sparsity based on a sampling strategy or it can be used as a masking matrix to record the observation state of the nodes.
| Notation | Definition |
| Underlying graph | |
| The set of nodes | |
| The set of edges | |
| N | The number of nodes |
| F | The number of features in each node |
| The vector of graph signals (node values) | |
| Identical matrix | |
| Adjacency matrix | |
| Degree matrix | |
| Laplacian matrix | |
| Normalized adjacency matrix | |
| Loss function |
III Graph Acquisition and Representation Techniques
@@
What differentiates GSP from conventional signal processing or standard multivariate algorithms is its integration of graph structures and topological relationships alongside signal data. This distinct approach leverages the underlying graph topology to enhance data representation and analysis in GSP. Consequently, the formulation of an appropriate graph topology when time-varying graph signals are present is crucial for representing data within GSP and graph machine learning frameworks. Before delving into the algorithmic aspects of TV-GSP, it is essential to examine key techniques for representing multivariate signals on topological structures.
III-A Conventional Static Graph Acquisition and Representations
The simplest and most common way of acquiring the graph topology is when the signals naturally contain or form a graph-like or network-like structure. Real-world data represented on the nodes can be constructed based on realistic relationships. A common example of naturally occurring graph structures can be found in traffic forecasting tasks: in the real world sensors positioned at various locations along highways that record data such as traffic speed or flow are perceived as nodes in a graph; the edges in the graph represent the connections between nodes, corresponding to the spatial distribution of sensors across the highways [15]. In the case of meteorological data collection, variables such as longitude and latitude from several weather stations are aggregated and analyzed using a Gaussian Kernel to construct a k-nearest-neighbor graph [2]. A graph depicting the strength of 5G signal reception is created as detailed in [13] by documenting signal strength across various sensors. A slightly more abstract category of graph generated from the real world is the social network. For example, users on social media are represented as nodes, and interactions between users create edges between nodes [16]. Another example is a citation network in which each node is a scientific paper and edges represent citations from one paper to another [17]. While many real-world datasets inherently exhibit graph structures, in other cases, the graph topology may not appear as straightforward. Specialized methods are employed to construct graphs that capture underlying relationships in the data.
Under the scenario where direct inference of the graph topology is not straightforward, a graph can still be formed using some techniques. In statistical graphical models, statistical measures between two data elements are used to define the pairwise relationship. For example, the covariance matrix and the precision matrix can be used to form a graph because covariance is a statistical measure of the deviations between two variables, meaning that measuring and properly thresholding this pairwise relationship measure gives us a sparse graph [18]. The Gaussian Graphical Model is a specific example of a Statistical Graphical Model that represents a conditional dependency structure in a graph: the nodes of the graph are the multivariate Gaussian distributed variables, and edges between these variables are the conditional dependencies or non-zero partial correlation coefficients [19, 20]. In GSP operations such as graph construction, transforms, sampling, prediction, and regularization can be realized from a probabilistic and statistical aspect using the Gaussian Graphical Model [21]. The graph Laplacian can be learned using a merging of the Gaussian prior and a smoothness assumption on the multivariate signal, which is equivalent to learning the underlying topological structure [11]. A multivariate signal can be assumed to be a sparse linear combination of atoms from a dictionary based on a polynomial of graph Laplacians that represents localized patterns from a sparse prior [22]. Another sparse prior assumption topology learning approach can be found in [23]: the Gaussian Graphical Model is used in a way that data is trained in a neural network to learn a function calculating the empirical covariance matrices to form graph structures.
Graph embedding techniques leverage adjacency and topological properties to project signals and features onto an embedding space to facilitate the analysis, visualization, and application in downstream GSP tasks. In the spectral domain, graph embeddings can be achieved by leveraging the eigenvectors of graph representations such as the adjacency matrix [24]. In the spatial domain, early approaches to generating the graph embedding are to use the random walk to propagate through the graph and then embed the data on the walk trajectories as words using approaches such as work2vec [25]. The DeepWalk generates embeddings by performing random walks on the graph, treating each node sequence like a sentence in natural language processing [26]. The node2vec extended DeepWalk further by introducing a flexible search mechanism that balances breadth-first and depth-first search strategies [27]. Recently, a more common method to generate graph embeddings is through GNNs such as GCNs. The spectral GCN is achieved by feeding the spectral graph convolution in (4) into a nonlinear activation function and replacing the filter with learnable filters as if the filters are parameters leaned from backpropagation [28]. The spatial GCN and the ChebNet are the spatial domain versions of GNNs with the spectral convolutions approximated by the Chebyshev polynomials as shown in (6) and let the neural network learn the polynomial coefficient instead of the spectral filter [8, 10]. The GraphSAGE exploits graph aggregations that sample and combine features from node neighborhoods, enabling an inductive framework where the model generalizes better to unseen nodes during inference [16]. The message-passing neural networks (MPNNs) further generalize the previous GNNs by introducing a flexible and localized method that computations are defined at the node level based on purely the node local adjacencies instead of defining all computations at once [29]. These graph embedding techniques establish a foundation for integrating GSP with machine learning tasks, including classification and clustering, by transforming graph signals along with the information in the graph topology into representations suitable for non-graph machine learning models.
III-B Dynamic Graph Acquisition and Representation
The pairwise relationships in graphs can be dynamic. Time-evolving processes over graphs can be captured by kernels formed by a preselected dictionary that fits the signal of interest where partial correlation, partial Granger causality, non-linear structural equations, and vector autoregressions techniques are used along with assumptions of the graph signals such as low-rank, sparsity, and smoothness [30, 31]. A dynamic graph can be formed from the correlated properties, which are previously seen in static Statistical Graphical Models, in time-varying graph signals by considering the space-time interactions in the underlying signal [32, 33, 34]. The topology inference on dynamic signals can be approached using a state-space formulation of a network process [35]. The Graphical LASSO is another instance of the Statistical Graphical Model method that estimates the inverse covariance matrix (precision matrix) with sparsity constraints enforced through an penalty. By combining Graphical LASSO with an ADMM-based algorithm, one can incorporate signals from neighboring time points and pairwise connections to learn the smoothly changing graphs [36]. Another smoothness-based graph topology learning task is to use the Sobolev smoothness as a measure to define a new smoothness function for time-varying graph signals [37]. The work of Natali et al. proposes to use the structural equation model to learn graph topologies in an online fashion and further proposes a combination of the Gaussian graphical model, the structural equation model, and the smoothness-based model into an all-in-one online topology learning framework [38, 39]. The causal relationships between data can be modeled using a Vector Autoregressive model, making it possible to infer a graph topology from the underlying time-series vector [40]. A specific example of this approach can be seen in a time-varying gene activation graph formed by thresholding a regression matrix of a time-varying gene regression that reflects gene activation patterns over time, leading to a parametric representation of the time-evolution and influence between pairs of data [41]. Given an underlying graph structure at a one-time instance, assuming that the dynamics of a time-varying graph only have some minor changes between time instances, a small perturbation analysis of the Laplacian matrix and its GFT can be done to model the time-evolution of the dynamic graph [42]. Topology inference of a dynamic topology from time-varying data can be tackled from another angle on an iterative prediction-correction approach on a Gaussian Graphical Model [43].
III-C Representing High-order Signals on Graphs
Despite the success of GSP, data representation in GSP is solely on the graph nodes, meaning that higher-order signals, such as signals on graph edges or signals on structures formed by edges, are not being represented. Examples of signals existing on the graph but not only the graph nodes can be found in the following literature: traffic data of the road network being recorded on the edges [44], gene interaction over time forms a set of time-varying graph edges [41], multi-agent systems that use time-varying edge data to model the agent interactions [45].
Higher-order structures such as hypergraphs, simplicial complexes, cell complexes, and multi-graphs enable the possibilities of defining graph representations beyond pairwise relationships [46]. Researchers have introduced the concept of Topological Signal Processing (TSP) as an alternative perspective to GSP [47]. In TSP, the Hodge Laplacian, a more generalized form of the graph Laplacian, is used to represent higher-order features beyond individual graph nodes [48]. The Hodge Laplacian can be used to represent simplicial complexes or cell complexes [47, 49]. When using simplicial complexes as building blocks, signals on graph nodes are treated as signals on the order simplex, graph edges are treated as signals on the order simplex, and triangles formed by edges are treated as signals on the order simplex, enabling the representation of signals on edges and beyond [47, 50]. Signal smoothing and denoising can be achieved on the graph edges by the simplicial complex representation [51]. Simplicial signals can also be embedded into simplicial complexes using methods similar to graph (node) embedding [52, 53]. Semi-supervised learning is applied to edge data using the Hodge Laplacians and flow conservation principles [44]. An analogous concept of spectral wavelets for simplicial complexes, termed Hodgelets, is introduced in [54]. Topological Slepians were proposed as signals concentrated on the topological domain and localized in the frequency domain [55]. Finite impulse response filters on simplicial complexes are discussed and further defined within convolution operations for simplicial complex signals [56, 57, 58]. A Chebyshev polynomial approximation for filters on simplicial complexes is also presented in [58] and [59]. These advancements underscore the important role of high-order representation in enhancing GSP for complex data.
IV Graph Signal Processing for Time-varying Graphs

### v2

# Signals on Graphs Over Time: Methodologies, Applications, and Challenges

@@ Abstract
Graphs, as effective representations of multidimensional, non-Euclidean, and irregularly structured data, have attracted significant research attention in recent years and are widely applied in diverse real-world domains such as social networks, traffic systems, and energy infrastructures. However, conventional graph-based techniques postulate that the graph is static in both signal and topology. Given the inherently dynamic nature of many real-world systems, research on Time-Varying Graphs (TVGs) has emerged as a critical area, offering the capability to capture essential dynamics of signal and topology changes over time. In the field of Graph Signal Processing (GSP), foundational concepts such as shifting, convolution, filtering, and transformation are adapted to address the challenges presented by time-varying signals on graphs. These techniques are further extended to enhance the implementation and interpretation of Graph Neural Networks (GNNs). Despite rapid progress, the relationships between GSP formulations and GNN architectures for TVGs remain fragmented. This review surveys recent advances in algorithms, applications, and benchmark datasets for TVGs, providing a unified perspective on how GSP principles inform the design and interpretation of graph neural networks for dynamic networks.
1 Introduction
In the present intelligent era, large-scale data collected from complex systems, such as urban transportation networks, meteorological observations, and brain connectivity, are increasingly represented as graphs and networks [171, 170, 70]. This representation enables the analysis of structured and interdependent data that traditional Euclidean models cannot adequately describe. Early studies in graph-based data analysis and Graph Signal Processing (GSP) typically assume that the underlying graph structure remains static, allowing classical signal processing concepts to be extended to graph domains, including graph shift, graph convolution, Graph Fourier Transform (GFT), Graph Wavelet Transform (GWT), and various graph filters such as low-pass, high-pass, and band-pass filters [167, 130, 60, 178]. However, many real-world systems exhibit inherently dynamic behavior in which both the graph signals and network topology evolve. For example, brain connectivity patterns vary across cognitive states [215], while transportation and environmental monitoring networks continuously change due to temporal fluctuations in demand and physical conditions [207]. These dynamics challenge the assumptions of static graph models and motivate the development of signal processing frameworks capable of handling time-varying graphs (TVGs).
Even though treating TVGs as multivariate time series provides a straightforward approach for data processing, this simplification ignores the interactions among observations that are encoded in the underlying graph topology. Such interactions are often essential for understanding the dynamics of many real-world systems. For example, in missing-not-at-random scenarios where observations associated with certain variables or entities are consistently unavailable, the missing entries may correspond to features that are systematically unobserved or unreliable [33]. In many real-world systems, these variables are not independent but exhibit strong dependencies. Exploiting these relationships allows missing values to be inferred from related observations, whereas approaches that ignore such dependencies struggle to produce reliable estimates. Moreover, conventional time-series models typically rely on independent or identically distributed assumptions, which fail to capture the relational dependencies naturally represented by graph connectivity. These limitations highlight the need for dedicated algorithms and models capable of processing signals on TVGs.
As an extension of GSP, time-varying Graph Signal Processing (TVGSP) has emerged to address dynamic graph data by modeling evolving signals through adaptive updating and spatiotemporal relationship estimation. Meanwhile, graph neural networks (GNNs), with their strong capability for nonlinear and complex representation learning, have become powerful tools for processing graph-structured data [86, 183]. Building on this framework, time-varying Graph Neural Networks (TVGNNs) have been developed to address dynamic graph scenarios through spatiotemporal learning and continual learning mechanisms [38, 185]. In both TVGSP and TVGNN approaches, a central challenge is the effective modeling of joint variations across spatial and temporal domains. TVGSP and TVGNN both focus on modeling and learning from graph-structured data that evolves over time, providing complementary tools for understanding and processing spatiotemporal graph signals. In particular, TVGSP provides theoretical foundations for understanding the evolution of signals on dynamic graphs, where insights into the behavior of time-varying data can inform the design and interpretation of TVGNN architectures.
In this review, we aim to systematically connect TVGSP and TVGNNs by examining their methodologies and analyzing their respective strengths and limitations. While studies have reviewed GSP [130, 95, 39, 74] and GNNs [27, 192, 223, 151, 25], existing surveys primarily focus on static graphs with time-invariant data. Several recent surveys have provided comprehensive overviews of TVGNNs from deep-learning perspectives, including spatial models and tasks for dynamic graphs [54, 109, 218], spatio-temporal GNN architectures [26, 156, 80], and dynamic graph representation learning frameworks [202]. These works offer valuable coverage of TVGNN taxonomies, datasets, and application trends. However, most previous studies examine only one paradigm in isolation, either GSP or GNN, thereby overlooking the fundamental connections between TVGSP and TVGNNs. In contrast, this review focuses on bridging TVGSP and dynamic GNN models, highlighting how GSP formulations, spectral interpretations, and filtering principles provide theoretical insights that inform, complement, and sometimes extend neural spatiotemporal architectures. As closely related approaches for modeling signals on TVGs, a unified review of TVGSP and TVGNNs can offer deeper insights into dynamic graph learning and help guide future research directions. The main contributions are summarized as follows:
Comprehensive framework: We provide a structured review of the fundamental concepts and defining characteristics of TVGSP and TVGNNs, establishing a cohesive framework that organizes key techniques and representative models.
Unified perspective: We present a unifying perspective that connects mainstream TVGNN approaches with TVGSP principles, clarifying their relationships and discussing advantages, limitations, trade-offs, and practical considerations in real-world TVG applications.
Future research directions: We discuss open challenges and emerging opportunities in processing TVGs, including new perspectives on discrete-time dynamic graphs (DTDG) and continuous-time dynamic graphs (CTDG) within the TVGSP framework, and highlight promising research directions.
@@ 1 Introduction
In Section 2, we begin with important definitions and background knowledge about TVGSP and TVGNNs. In Section A, we focus on the acquisition and representation of TVGs. Sections 3 and 4 provide a comprehensive review and discussion of the existing approaches for processing TVGs, including both signal processing and neural network methods. Section 5 summarizes the commonly used datasets, the corresponding application, and the evaluation metrics. Finally, Section 7 concludes and indicates the future direction of the TVGSP.
2 Preliminaries
Graphs effectively encapsulate complex interactions within multivariate systems, where pairwise relationships between variables are encoded as edges. The nodes, edges, and the data associated with them can be dynamically updated to reflect the intrinsic behavior and evolving nature of the underlying relationships. Graph representations emerge and capture the structure and interdependencies of the data.
2.1 Definition of Time-Varying Graphs and Signals
@@
In conventional GSP that operates on static data, the main research focus is placed on a static graph , which is defined by a set of nodes and a set of edges that represent the connections between nodes. In this review, the graph is assumed undirected unless otherwise specified, and can be either weighted or unweighted. A graph signal is defined as a multi-dimensional vector, recording the value of each node in . We further denote the number of nodes with . The adjacency matrix of comprises the topological information by storing the state of connectivity between nodes. To be specific, the entry of is the edge weight between nodes and . If is unweighted, indicates there is an edge between nodes and . Conversely, means there is no connection between and . For an undirected graph, the adjacency matrix is always symmetric. If is undirected and unweighted, the number of edges connected to a node is defined as the node degree , which can be represented by a diagonal matrix called the degree matrix . In the weighted case, the degree of a node is the sum of all edge weights connected to it, rather than simply counting the number of edges.
Since real-world data often evolves over time, meaning that its graph representations, including node values, edge weights, and even graph topology, could be dynamic, applying static techniques independently at each time instance is inadequate, as it neglects the temporal dependencies inherent in dynamic systems. Consequently, a variety of TVG algorithms have been developed to address different forms of graph dynamics. In this review, following previous works [80, 42, 218], we categorized TVGs into three classes from the evolutionary perspective: the Spatiotemporal Graphs (STGs), the Discrete-Time Dynamic Graphs (DTDGs), and the Continuous-Time Dynamic Graphs (CTDGs) as illustrated in Fig. 1.
Given multi-dimensional data sequence, , where is the number of vertices and is the number of time points. A TVG is noted by , where is the vertex set, and respectively denote the edge set and adjacency matrix at time t. In most TVG modeling conventions, the vertex set is assumed to remain constant, while the edge set may evolve over time. Accordingly, the associated adjacency matrix , edge-based operators (e.g., ), and edge-dependent quantities change consistently with , whereas node insertion or deletion is typically treated as a separate modeling regime.
In STGs, the node signals vary over time, while other graph elements, such as the topology or the adjacency matrix , remain static. STGs are also known as the time-vertex graphs, a terminology commonly used in the GSP literature [52, 85], and STGs are more frequently seen in GNN-related research [107, 80]. In GSP, time-varying signals on STGs can be modelled using product graphs. Product graphs capture not only correlations between nodes but also correlations across time [79]. In both DTDGs and CTDGs, the dynamic behaviors are shown in not only the signals on the graph nodes but also in the graph topology itself . Specifically, in DTDGs, the dynamic graphs are represented by a series of snapshots that can be the graph adjacency matrices and corresponding node feature vectors over time. Both STGs and DTDGs are extensively used, as time-varying dynamics are straightforwardly represented as snapshots. By contrast, the CTDGs have emerged in the field of TVGNNs and have yet to be explored in TVGSP. In CTDGs, the TVGs are represented by a more complicated representation in which an initial graph is first defined, and a sequence of graph evolution events defines the time-varying dynamics over time.
2.2 Graph Signal Processing and Graph Neural Networks
@@
The graph Laplacian matrix , which combines information from and , is defined as . The GFT is based on the eigenvector decomposition
where is the orthonormal eigenvector matrix and is the diagonal matrix of eigenvalues diag. Notice that for STGs, the GFT in (1) is directly applicable to all time points since the Laplacian is fixed over time. For DTDG and CTDG settings, the same definition still holds, but the GFT is applied to each graph realization, meaning (and its eigendecomposition) is recalculated whenever the topology changes.
@@
The GFT transforms a graph signal from the spatial domain to the spectral domain by projecting onto :
@@
The GFT of the TVG signal (hourly temperature) in Fig. 3(a) is shown in Fig. 3(b). Spectral-domain operations can be performed similarly to those in the classical Fourier transform by defining a filter , which is then applied using the convolution property of the Fourier transform. A basic yet complete graph convolution procedure to apply the filter to in the spectral domain and transform it back into the spatial domain:
In GSP, concepts like low-pass, high-pass, band-pass, and band-reject filters can be analogously implemented through filter design [178]. Wavelets on graphs can be constructed by applying a kernel function to the decomposition of the graph Laplacian in (1) to achieve localized representations [60]. A graph signal is spectrally sparse when it is bandlimited in the frequency domain. For example, a bandlimiting filter is defined based on a frequency set , in which , with if and 0 otherwise [31]. A (spectral) GCN layer can be obtained by feeding (4) into a non-linear activation :
where is an additional bias term to implement, and will now be learned from the back-propagation as learnable parameters instead of being predefined.
The spectral graph convolution can be approximated into a spatial graph convolution by using Chebyshev polynomial approximation for a series of Chebyshev polynomials defined over the interval [168]:
@@
A graph signal with a reduced number of nodes, sampled or masked according to a node sampling set , is sparse in the spatial domain. The matrix represents the sampling or masking matrix, containing only non-zero diagonal elements given by . The matrix can be designed to introduce spatial sparsity based on a sampling strategy, or it can be used as a masking matrix to record the observation state of the nodes. Similar to how the (spectral) GCN is defined in (5), a (spatial) GCN can be defined using (7):
where parameters are the learnable GCN parameters and is again the bias term. It should be emphasized that equations (5) and (8) establish the basic but direct relationship between GSP and GNNs, making GNNs interpretable using GSP techniques. GSP interpretations of more advanced GNNs can be achieved through appropriate redefinition of trainable parameters, modification of core components, and integration of machine learning techniques from (5) and (8).
@@
What differentiates GSP from conventional signal processing or standard multivariate algorithms is its integration of graph structures and topological relationships alongside signal data. This distinct approach leverages the underlying graph topology to enhance data representation and analysis in GSP [118]; thus, methods that use conventional spectral transformations, such as DFT, will be omitted.
3 Signal Processing for Time-varying Graphs